# On the Scalability of GNNs for Molecular Graphs

### v3

@@ 1 Introduction
Figure 1: Summary of our GNN scaling hypotheses studied in the present work. The baseline model is presented in dark grey, followed by different scaling hypotheses illustrated in lighter colors. We analyze the scaling behavior of message-passing networks, graph Transformer and hybrid architectures with respect to the increasing scale of width, depth, number of molecules, number of labels, and diversity of datasets.
Recent successes in language modelling (OpenAI, 2023; Touvron et al., 2023) and image generation (Ramesh et al., 2021; Rombach et al., 2022) are driven by the increasing amount of training data and computational resources. Across different domains, practitioners have observed a direct relationship between model parameter count and performance on novel tasks (Kaplan et al., 2020). In natural language processing, large Transformer-based models have demonstrated impressive generalization capabilities utilizing a causal autoregressive objective (Radford et al., 2019). In the meantime, image generation has undergone incredible leaps with large models trained utilizing pixel level unsupervised objectives.
While data power law scaling behavior has been tremendously beneficial in language and image domains, its practical impact on molecular reasoning and drug discovery has remained limited. This is a direct consequence of complex scientific tasks requiring reasoning regarding the underlying structure of the data (Bubeck et al., 2023). In the past, molecular property prediction approaches have made use of graph-based methods, as these allow us to reason about the structure and interaction of different components of a molecule. Molecules are naturally represented as graphs, where the nodes represent atoms and edges correspond to covalent bonds between the atoms.
Graph Neural Networks (GNNs) have emerged as a promising way of learning molecular representations (Liu et al., 2023a; Galkin et al., 2023). GNN architectures are equipped with the flexibility of learning molecular structure while building general, powerful representations over graphs utilizing backpropagation. These representations have been utilized in various paradigms such as reasoning about drug properties (Stärk et al., 2022a), target binding interactions (Stärk et al., 2022b), retrosynthesis of reagents (Liu et al., 2020), ligand-based docking (Corso et al., 2023) and in-silico experiment design (Wang et al., 2023b).
@@
Learning molecular properties with GNNs presents its own set of unique challenges. First, multiple different GNN architectures are being actively researched. These include convolution (Kipf & Welling, 2017), message passing architectures (Beaini et al., 2021), graph Transformers (Ying et al., 2021b) and hybrid architectures (Rampášek et al., 2022; Masters et al., 2022). These approaches have shown recent progress, but their applicability to practical applications remains an open question (Rong et al., 2020).
Second, commonly used self-supervised training techniques do not transfer well when applied to molecular graphs; e.g., retrieving masked bonds and atoms is not informative. This is primarily due to large data requirements and the fact that graphs are limited in capturing domain-specific aspects such as chemical interactions and biological compositions (Liu et al., 2022). Other methods such as GPSE (Liu et al., 2023b) solely learn the graph structure, thus providing a better positional encoding for another GNN.
Lastly, public datasets have insufficient high-quality data for effective GNN architecture training. While recent attempts have been made to expand open-source datasets (Beaini et al., 2024), their extensions towards multi-task settings remain an open question.
We aim to address these limitations and provide a concrete understanding of the required data and architectures to build foundational models for molecular graphs. Specifically, we want to answer the question: How do graph-based architectures scale in multi-task settings of large molecular graphs?
@@
Our work aims to provide a proper understanding of how different GNN architectures scale for molecular GNNs and how it affects their performance in various settings. Our main contributions are as follows:
@@
  • We show that our largest 1 billion parameter models continue to scale with constant gains in molecular property prediction. To the best of our knowledge, this is the first work to demonstrate the continuous scaling behavior of molecular GNNs.
@@
  • We provide an in-depth analysis of scaling trends across different probing and finetuning strategies. Specifically, we observe that model width and number of labels are the most important factors driving finetuning performance.
@@
2 Related Work
Foundation Models for Molecules. Recent work has relied on foundation models as a generalist class of models for sequential modelling (Yuan, 2023; Liu et al., 2023a) as well as knowledge retrieval (Galkin et al., 2023). Within molecular drug discovery, recent works rely on structured models of ligands (Moret et al., 2020). Luo et al. (2022) and Moret et al. (2023) study a general model for protein synthesis. Rao et al. (2020) construct a self-attention driven architecture for contact prediction. Madani et al. (2023) learn to generate a family of functional proteins. Nijkamp et al. (2023) present a class of protein-pretrained language models. Similarly, Méndez-Lucio et al. (2022) study binding interactions between different assays at the molecule-molecule interaction level.
While many models focus on the design of molecules, a recent class of methods has also focused on properties of molecules (Beaini et al., 2024). Cui et al. (2023) and Luo et al. (2023) study a general architecture for predicting similar properties across different molecules such as solubility and viscosity that in a way generalizes across molecular domains with a limited set of samples. Unsal et al. (2022) learn to predict functional properties of proteins by exploiting learned structure. Our study explores similar molecular tasks for property prediction.
Architecture Design. Recent methods in graph architecture design focus on attending to structural information across nodes (Müller et al., 2023). Of specific interest are graph Transformer networks which extract node as well as edge information by composing sequential attention modules over graph readouts (Yun et al., 2019). In parallel, graph attention networks model attention weights across edges of a graph (Veličković et al., 2017).
While attention mechanisms have demonstrated modern progress, traditional architectures such as neural message passing (Gilmer et al., 2017) and 3D infomax (Stärk et al., 2022a) hold the promise of simplicity and expressivity in modelling molecular graphs. On one hand, message passing provides a rich and expressive framework for constructing representations. On the other hand, the provable lower bound of infomax results in strong convergence guarantees. Godwin et al. (2021) study regularization based on noisy nodes for the task of molecular property prediction. Provision of noise imputation in node-level features leads to simple and expressive method for tackling sparse molecular graphs. Graph bootstrapping (Thakoor et al., 2021) allows prior architectures to scale up to larger and complex graphs for representation learning. Our exploration of different architectures is aligned with the aforesaid works in literature, and with recent trends towards Transformers in related machine learning fields.
Scaling Laws. Recent work in model scaling has demonstrated that performant models follow a power law relationship between their parameter sizes and performance on new data samples (Kaplan et al., 2020). Additionally, this relationship holds during the finetuning stage (Hernandez et al., 2021), thus indicating a strong reliance on model parameters. Bahri et al. (2021) explain this power law fit by observing learning as moving on a smooth data manifold. Frantar et al. (2023) study the power law fit for sparsely connected models capable of downstream generalization. Notably, sparsely connected foundation models reach a computational bottleneck as a result of different sparsity structures impacting hardware usage. Aghajanyan et al. (2023) study the power law fit for mixed modality generative models, indicating that the scaling behavior is modality agnostic across various datasets. The result hints at the generality of scaling across different domains and applications.
Caballero et al. (2022) extend the study of scaling laws towards different training regimes (such as finetuning, downstream transfer and inference) as well as different problem domains (vision, language, audio, diffusion, generative modelling, contrastive learning, and reinforcement learning). The resulting functional form results in extrapolations which are empirically accurate. Diaz & Madaio (2023) present a tangential result demonstrating that dataset entities may not necessarily scale with the growing amount of parameters and computational requirements, likely due to models leaving out essential data samples. Cherti et al. (2023) study the reproducibility of scaling laws for contrastive learning scenarios at the intersection of language and vision.
While recent work on the scaling behavior of graph networks and other structural inductive biases remains absent, a few notable works aim at studying the behavior of scaling graph sizes. Grindrod (2022) evaluates the power law fit of growing graph sizes utilizing random walk sampling and exploration. Similarly, Pražnikar (2021) studies the scaling behavior of macromolecule proteins with respect to their mean node degree. Our exploration of scaling behaviors in graph networks is motivated by the aforesaid directions.
@@ 3.1 Graph Neural Networks
Our problem setting consists of graphs of the form $\mathcal{G}=(\mathcal{V},\mathcal{E})$ where $\mathcal{V}$ denotes the set of nodes and $\mathcal{E}$ denotes the set of edges. Each node $i\in\mathcal{V}$ indicates the atom and each edge $(u,v)\in\mathcal{E}$ denotes a chemical bond between two atoms $u,v$. Total number of atoms in the molecule are $N=|\mathcal{V}|$ while total number of edges are $M=|\mathcal{E}|$. Node features in layer $l$ are denoted by $\mathbf{x}_{i}^{l}\in\mathbb{R}^{d}$ and are concatenated into an $N\times d$ representation matrix $\mathbf{X}^{l}=[\mathbf{x}_{1}^{l};\mathbf{x}_{2}^{l};...\mathbf{x}_{N}^{l}]^{\top}$. Edge features $e_{uv}^{l}\in\mathbb{R}^{d}$ are concatenated into the edge feature matrix $E^{l}=[e_{uv}^{l}:(u,v)\in\mathcal{E}]^{\top}$.
@@ 3.2 Scaling Laws
We denote the parameters of a model as $\theta$ with the total number of trainable parameters being $|\theta|$. We consider a training dataset $\mathcal{D}$ consisting of labeled data samples $(\mathcal{G},y)\in\mathcal{D}$. Here, $\mathcal{G}$ indicates the input graph and $y\in\mathbb{R}^{N}$ denotes the categorical or continuous label. Total size of the dataset is denoted as $|\mathcal{D}|$. Given the study of large foundational models, we note that $|\theta|$ is large in size and $\theta$ lies on a high dimensional manifold such that $\theta\in\mathbb{R}^{B}$ where $B>>|\mathcal{D}|$. Recent work has shown that increasing the size of dataset $|\mathcal{D}|$ or the number of trainable parameters $|\theta|$ has a direct power law relationship on the loss function $L_{\theta}(|\mathcal{D}|)$ (Kaplan et al., 2020). Mathematically, we have the following,
$L_{\theta}(|\mathcal{D}|)\propto(|\theta_{C}|/|\theta|)^{\alpha}\vspace{-0.30cm}$ (1)
@@
$L_{\theta}(|\mathcal{D}|)\propto(|\mathcal{D}_{C}|/|\mathcal{D}|)^{\beta}\vspace{-0.30cm}$ (2)
Equation 2 describes the power-law relationship between the size of dataset and loss obtained when training the model on $\mathcal{D}$. Here, $|\mathcal{D_{C}}|$ denotes the critical size of the dataset and $\beta\in\mathbb{R}$ is a scalar constant.
@@ 4 How Do Molecular GNNs Scale?
Our study aims to answer the question: How do molecular GNNs scale? We begin by studying GNNs in the multi-task supervised pretraining setup. Since our analysis consists of multiple tasks on a large scale, we utilize the Graphium library (Beaini et al., 2024). Due to the absence of a unified consensus on the best architecture for molecular GNNs, we focus our efforts on three specific models. We select MPNN++ (Masters et al., 2022) which improves quantum prediction over the MPNN (Gilmer et al., 2017), Graph Transformers (Müller et al., 2023), and Hybrid GPS++ (Masters et al., 2022) along with the use of positional encodings. Finally, we evaluate our models on a range of public benchmarks with 38 datasets from TDC (Huang et al., 2021), Polaris [^1] PolarisHub: https://polarishub.io/, and MoleculeNet (Wu et al., 2018). We evaluate our models in both finetuning and probing (fingerprinting) settings.
@@ 4.1 Datasets
We study the scaling behavior of GNNs on the LargeMix dataset mixture (Beaini et al., 2024). These datasets cover different types of molecules exhibiting variable properties. Thus, the training is done in a multi-task setting consisting of thousands of labels. This is a challenging approach towards learning representations with GNNs, especially as some labels are very imbalanced and sparse.
@@
  • PCBA_1328 is a dataset of 1.6M molecules with 1,328 binary classification labels. Each label corresponds to the activity tags of the molecules in a bioassay reported on PubChem (Kim et al., 2023).
  • PCQM4M_G25 and PCQM4M_N4 are two datasets of 3.8M molecules with 25 graph-level labels and 4 node-level labels. Labels are obtained using Density Functional Theory (DFT) simulations, a highly accurate quantum simulation method (Saal et al., 2013).
@@ 4.2 Finetuning and Probing Benchmarks
TDC. Therapeutics Data Commons (Huang et al., 2021) is one of the common benchmarks for drug discovery. Our study focuses on 22 ADMET (Absorption, Distribution, Metabolism, Excretion and Toxicity) tasks. While TDC serves as the bedrock for open-source drug discovery evaluation, we note that it suffers from data collection and processing biases across dissimilar molecules (Walters, 2023).
@@
MoleculeNet. This is a benchmark dataset for molecular machine learning that is built upon public datasets (Wu et al., 2018). It consists of various datasets covering different levels of molecular properties spanning from properties at the molecular level to broader impacts on the human body. There are different categories of properties including quantum mechanics, physical chemistry, biophysics, and physiology. We investigate four datasets that are commonly used in similar studies such as (Liu et al., 2021; Xu et al., 2021b; Liu et al., 2023b).
@@ 4.3 Architectures
MPNN++. This is a variation of the neural message passing architecture with edge and global features (Gilmer et al., 2017; Battaglia et al., 2018; Bronstein et al., 2021). Choosing the MPNN++ allows us to maximize architecture expressivity while minimizing the risk of overfitting on larger datasets (Masters et al., 2022). Each MPNN block makes use of sequential Dropout (Srivastava et al., 2014), MLP and LayerNorm (Ba et al., 2016) modules followed by a skip connection (He et al., 2016b; Srivastava et al., 2015) across node and edge features:
$\bar{E}^{l},\mathbf{X}^{l}$ $=\texttt{Dropout}(\texttt{MLP}([\mathbf{X}^{l}|E^{l}]))$
$\mathbf{X}^{l}$ $=\texttt{LayerNorm}(\texttt{Dropout}(\mathbf{X}^{l}))+\mathbf{X}^{l}$
$E^{l+1}$ $=\bar{E}^{l}+E^{l}$
GPS++. This is a hybrid model leveraging the MPNN++ inductive bias while providing the flexibility of self-attention-based modules (Ying et al., 2021a) to allow for a rich feature extraction scheme across nodes and edges, and was empirically proven very successful (Masters et al., 2022). Here, the standard self-attention weights are biased by a structural prior $\mathcal{B}$ from the input graph. Mathematically, the GPS++ module carries out the following computation:
[…]
$\mathbf{Z}^{l}$ $=\texttt{BiasedAttn}(\mathbf{X}^{l+1},\mathcal{B})$
$\mathbf{X}^{l+1}$ $=\texttt{MLP}(\mathbf{X}^{l+1}+\mathbf{Z}^{l})$
Transformer. This is an architecture identical to the GPS++, but without the MPNN++ module nor the concatenation. Instead, it relies solely on the PSEs for structural bias.
PSEs. These are important design choices when training GNN architectures (Rampášek et al., 2022; Liu et al., 2023b), as they allow each node to understand its position and surroundings within a graph. This is essential for any graph Transformer, but it was also shown to improve the expressivity of molecular GNNs. Specifically, we use three PSE schemes. First, we use random-walk diagonals (Dwivedi et al., 2021) to allow one to decouple structural and positional representations. Learned positional encodings are used to tackle isomorphic nodes. Second, we use Laplacian eigenvectors (Beaini et al., 2021) as these form an expressive way to encode node geometries and positions. Laplacian eigenvectors provide strong theoretical guarantees with respect to the expressivity of the Weisfeiler-Lehman test, a useful insight when evaluating GNNs at scale. Last, we use the Laplacian eigenvalues (Kreuzer et al., 2021) as a suitable PSE scheme to fully leverage the Laplacian spectrum. Additionally, they provide global structural information about the graph.

### v4

@@ 1 Introduction
Figure 1: Summary of our GNN scaling hypotheses studied in the present work. The baseline model is presented in dark grey, followed by different scaling hypotheses illustrated in lighter colors. We analyze the scaling behavior of message-passing networks, graph Transformers and hybrid architectures with respect to the increasing scale of width, depth, number of molecules, number of labels, and diversity of datasets.
Recent successes in language modelling [47, 65] and image generation [52, 55] are driven by the increasing amount of training data and computational resources. Across different domains, practitioners have observed a direct relationship between model parameter count and performance on novel tasks [28]. In natural language processing, large Transformer-based models have demonstrated impressive generalization capabilities utilizing a causal autoregressive objective [51]. In the meantime, image generation has undergone incredible leaps with large models trained utilizing pixel level unsupervised objectives.
While data power law scaling behavior has been tremendously beneficial in language and image domains, its practical impact on molecular reasoning and drug discovery has remained limited. This is a direct consequence of complex scientific tasks requiring reasoning regarding the underlying structure of the data [10]. In the past, molecular property prediction approaches have made use of graph-based methods, as these allow us to reason about the structure and interaction of different components of a molecule. Molecules are naturally represented as graphs, where the nodes represent atoms and edges correspond to covalent bonds between the atoms.
Graph Neural Networks (GNNs) have emerged as a promising way of learning molecular representations [33, 17]. GNN architectures are equipped with the flexibility of learning molecular structure while building general, powerful representations over graphs utilizing backpropagation. These representations have been utilized in various paradigms such as reasoning about drug properties [61], target binding interactions [62], retrosynthesis of reagents [32], ligand-based docking [13] and in-silico experiments [70].
@@
Learning molecular properties with GNNs presents its own set of unique challenges. First, multiple different GNN architectures are being actively researched. These include graph-convolutions [30], message passing architectures [6], graph Transformers [80] and hybrid architectures [53, 39]. These approaches have shown recent progress, but their applicability to practical applications remains an open question [56].
Second, commonly used self-supervised training techniques do not transfer well when applied to molecular graphs; e.g., retrieving masked bonds and atoms is not informative. This is primarily due to large data requirements and the fact that graphs are limited in capturing domain-specific aspects such as chemical interactions and biological compositions [36]. Other methods such as GPSE [34] solely learn the graph structure, thus providing a better positional encoding for another GNN.
Lastly, public datasets have insufficient high-quality data for effective GNN architecture training. While recent attempts have been made to expand open-source datasets [7], their extensions towards multi-task settings remain an open question.
We aim to address these limitations and provide a concrete understanding of the required data and architectures to build foundational models for molecular graphs. Specifically, we want to answer the question: How do graph-based architectures scale in supervised multi-task training on molecular graphs?
@@
Our work aims to provide a proper understanding of how different GNN architectures scale for molecular GNNs and its affects on performance in various settings. Our main contributions are:
@@
  • We show that our largest 3 billion parameter models continue to scale with constant gains in molecular property prediction. To the best of our knowledge, this is the first work to demonstrate the continuous scaling behavior of molecular GNNs.
@@
  • We provide an in-depth analysis of scaling trends across different probing and finetuning strategies. Specifically, we observe that model width and number (and quality) of labels are the most important factors driving finetuning performance.
@@ 2.1 Graph Neural Networks
Our problem setting consists of graphs of the form $\mathcal{G}=(\mathcal{V},\mathcal{E})$ where $\mathcal{V}$ denotes the set of nodes and $\mathcal{E}$ denotes the set of edges. Each node $i\in\mathcal{V}$ indicates the atom and each edge $(u,v)\in\mathcal{E}$ denotes a chemical bond between two atoms $u,v$. Total number of atoms in the molecule are $N=|\mathcal{V}|$ while total number of edges are $M=|\mathcal{E}|$. Node features in layer $l$ are denoted by $\mathbf{x}_{i}^{l}\in\mathbb{R}^{d_{n}}$ and are concatenated into an $N\times d_{n}$ representation matrix $\mathbf{X}^{l}=[\mathbf{x}_{1}^{l};\mathbf{x}_{2}^{l};...\mathbf{x}_{N}^{l}]^{\top}$. Edge features $e_{uv}^{l}\in\mathbb{R}^{d_{e}}$ are concatenated into the $M\times d_{e}$ edge feature matrix $E^{l}=[e_{uv}^{l}:(u,v)\in\mathcal{E}]^{\top}$.
@@ 2.2 Scaling Laws
We denote the parameters of a model as $\theta$ with the total number of trainable parameters being $|\theta|$. We consider a training dataset $\mathcal{D}$ consisting of labeled data samples $(\mathcal{G},y)\in\mathcal{D}$. Here, $\mathcal{G}$ indicates the input graph and $y\in\mathbb{R}^{N}$ denotes the categorical or continuous label. Total size of the dataset is denoted as $|\mathcal{D}|$. Given the study of large foundational models, we note that $|\theta|$ is large in size and $\theta$ lies on a high dimensional manifold such that $\theta\in\mathbb{R}^{B}$ where $B>>|\mathcal{D}|$. Recent work has shown that increasing the size of dataset $|\mathcal{D}|$ or the number of trainable parameters $|\theta|$ has a direct power law relationship on the loss function $L_{\theta}(|\mathcal{D}|)$ [28]. Mathematically, we have the following,
$L_{\theta}(|\mathcal{D}|)\propto(|\theta_{C}|/|\theta|)^{\alpha}$ (1)
@@
$L_{\theta}(|\mathcal{D}|)\propto(|\mathcal{D}_{C}|/|\mathcal{D}|)^{\beta}$ (2)
Equation 2 describes the power-law relationship between the size of dataset and loss when training the model on $\mathcal{D}$. Here, $|\mathcal{D_{C}}|$ denotes the critical size of the dataset and $\beta\in\mathbb{R}$ is a scalar constant.
@@ 3 How Do Molecular GNNs Scale?
Our study aims to answer the question: How do molecular GNNs scale? We begin by studying GNNs in the multi-task supervised pretraining setup. Since our analysis consists of multiple tasks on a large scale, we utilize the Graphium library [7]. Due to the absence of a unified consensus on the best architecture for molecular GNNs, we focus our efforts on three specific models. We select MPNN++ [39] which improves quantum prediction over the MPNN [18], Graph Transformers [44], and Hybrid GPS++ [39] along with the use of positional encodings. Finally, we evaluate our models on a range of public benchmarks with 38 datasets from TDC [26], Polaris [^1] PolarisHub: https://polarishub.io/, and MoleculeNet [71]. We evaluate our models in both finetuning and probing (fingerprinting) settings.
@@ 3.1 Datasets
We study the scaling behavior of GNNs on the LargeMix dataset mixture [7]. These datasets cover different types of molecules exhibiting variable properties. Thus, the training is done in a multi-task setting consisting of thousands of labels. This is a challenging approach towards learning representations with GNNs, especially as some labels are very imbalanced and sparse.
@@
  • PCBA_1328 is a dataset of 1.6M molecules with 1,328 binary classification labels. Each label corresponds to the activity tags of the molecules in a bioassay reported on PubChem [29].
  • PCQM4M_G25 and PCQM4M_N4 are two datasets of 3.8M molecules with 25 graph-level labels and 4 node-level labels. Labels are obtained using density functional theory (DFT) simulations, a highly accurate quantum simulation method [57].
@@ 3.2 Finetuning and Probing Benchmarks
TDC. Therapeutics Data Commons [26] is one of the common benchmarks for drug discovery. Our study focuses on 22 ADMET (Absorption, Distribution, Metabolism, Excretion and Toxicity) tasks. While TDC serves as the bedrock for open-source drug discovery evaluation, we note that it suffers from data collection and processing biases across dissimilar molecules [68].
@@
MoleculeNet. This is a benchmark dataset for molecular machine learning that is built upon public datasets [71]. It consists of various datasets covering different levels of molecular properties spanning from properties at the molecular level to broader impacts on the human body. There are different categories of properties including quantum mechanics, physical chemistry, biophysics, and physiology. We investigate 4 datasets that are commonly used in similar studies such as [35, 74, 34].
@@ 3.3 Architectures
MPNN++. This is a variation of the neural message passing architecture with edge and global features [18, 5, 9]. Choosing the MPNN++ allows us to maximize architecture expressivity while minimizing the risk of overfitting on larger datasets [39]. Each MPNN block makes use of sequential Dropout [59], MLP and LayerNorm [3] modules followed by a skip connection [23, 60] across node and edge features:
$\bar{E}^{l},\mathbf{X}^{l}=\texttt{Dropout}(\texttt{MLP}([\mathbf{X}^{l}|E^{l}]));\quad\mathbf{X}^{l}=\texttt{LayerNorm}(\texttt{Dropout}(\mathbf{X}^{l}))+\mathbf{X}^{l};\quad E^{l+1}=\bar{E}^{l}+E^{l}$
GPS++. This is a hybrid model leveraging the MPNN++ inductive bias while providing the flexibility of self-attention-based modules [79] to allow for a rich feature extraction scheme across nodes and edges, and was empirically proven very successful [39]. Here, the standard self-attention weights are biased by a structural prior $\mathcal{B}$ from the input graph. Mathematically, the GPS++ module carries out the following computation:
$\mathbf{X}^{l+1},E^{l+1}=\texttt{MPNN++}(\mathbf{X}^{l},E^{l});\quad\mathbf{Z}^{l}=\texttt{BiasedAttn}(\mathbf{X}^{l+1},\mathcal{B});\quad\mathbf{X}^{l+1}=\texttt{MLP}(\mathbf{X}^{l+1}+\mathbf{Z}^{l})$
Transformer. This model is identical to GPS++, but without the MPNN++ module and concatenation. Instead, it relies solely on positional and structural encodings (PSEs) for structural bias.
PSEs. These are important design choices when training GNN architectures [53, 34], as they allow each node to understand its position and surroundings within a graph. This is essential for any graph Transformer, but it was also shown to improve the expressivity of molecular GNNs. Specifically, we use three PSE schemes. First, we use random-walk diagonals [15] to allow one to decouple structural and positional representations. Learned positional encodings are used to tackle isomorphic nodes. Second, we use Laplacian eigenvectors [6] as these form an expressive way to encode node geometries and positions. Laplacian eigenvectors provide strong theoretical guarantees with respect to the expressivity of the Weisfeiler-Lehman test, a useful insight when evaluating GNNs at scale. Last, we use the Laplacian eigenvalues [31] as a suitable PSE scheme to fully leverage the Laplacian spectrum. Additionally, they provide global structural information about the graph.